# Towards AI Urban Planner in the Age of GenAI, LLMs, and Agentic AI

### v1

@@ Abstract
Generative AI, large language models, and agentic AI have emerged separately of urban planning. However, the convergence between AI and urban planning presents an interesting opportunity towards AI urban planners. This paper conceptualizes urban planning as a generative AI task, where AI synthesizes land-use configurations under geospatial, social, and human-centric constraints. We survey how generative AI approaches, including VAEs, GANs, transformers, and diffusion models, reshape urban design. We further identify critical gaps: 1) limited research on integrating urban theory guidance, 2) limited research of AI urban planning over multiple spatial resolutions or angularities, 3) limited research on augmenting urban design knowledge from data, and 4) limited research on addressing real-world interactions. To address these limitations, we outline future research directions in theory-guided generation, digital twins, and human-machine co-design, calling for a new synthesis of generative intelligence and participatory urbanism.
@@ I Introduction and Motivation
II Rethinking Urban Planning in the Era of GenAI, LLMs, and Agentic AI
@@
II-A Urban Planning as Generative AI
GenAI, particularly models based on Generative Adversarial Networks (GANs), Variational Autoencoders (VAEs), autoregressive models, and diffusion processes, serves as the algorithmic backbones of automated urban layout synthesis. For instance, GANs utilize a generator to learn the distribution of well-planned spatial configurations, while a discriminator distinguishes between high-quality and poor-quality plans. An adversarial loss function orchestrates a minimax game to optimize both networks. VAEs encode contextual urban data of a target region to plan into a latent embedding and decode it into spatial configurations by minimizing reconstruction loss. Autoregressive Models employ sequential architectures (e.g., Transformers) to generate spatial configurations step-by-step, where each step is conditioned on the prior outputs and trained via maximum likelihood estimation. Diffusion Models learn data distributions through a forward noise-injection process and a reverse denoising process, enabling guided sampling from learned distributions to produce context-aware or goal-directed planning outcomes.
@@ II-A Urban Planning as Generative AI
II-B Recent Studies in GenAI for Urban Planning
@@
Inspired by the new abilities of GenAI, many domaine experts have leveraged these GenAI algorithms to reformulate urban planning as a generative AI task, where a spatial configuration plan can be seen as an image or a picture or a matrix or a tensor, and a generative model aims to generate the optimal image (or tensor) [1]. The study in [2] models urban planning as land-use configuration generation over attributed spatial graphs enriched with geographic context, where configurations are represented as multi-channel tensors encoding Point of Interest (POI) distributions by location and function. A graph autoencoder learns regional embeddings, which are input to a GAN that evaluates generated configurations using indicators such as willingness to pay, human activity diversity, and social interaction. While innovative, this model fails to capture spatial hierarchies and suffers from training instability. The study in [3] conditions land-use generation on spatial context and human textual guidance using a deep conditional variational autoencoder, producing zone maps and POI layouts guided by planner instructions and contextual data such as geography and mobility. Although effective in incorporating planning knowledge, it lacks explicit modeling of hierarchical urban structures. Addressing this, the study in [4, 1] proposes a hierarchical conditional transformer that first generates zone-level functions and then refines them into grid-level plans. The study in [5] models planning as a two-level hierarchical reinforcement learning (RL) task: an actor-critic network governs zone planning, while a Deep Q-Network handles POI-level design. Similarly, a related study [6] treats cities as graphs and applies GNN-based RL for sequential planning decisions.
II-C Limitations of Current AI Urban Planner Research
Despite advances in AI for urban planning, current approaches face critical limitations across five foundational dimensions: theory, methodology, data, computing, and applications. 1) Theory-wise, most AI models reduce planning to spatial optimization or land-use generation, thus, ignore core urban concepts such as zoning, resilience, social equity, and participatory governance, and often ignore the political and normative nature of planning decisions—undermining interpretability and legitimacy. 2) Method-wise, models [2, 6, 5, 4, 1, 3, 2] rely on single-step or loosely staged generation conditioned on static inputs that limit creativity and adaptivity, fail to capture multi-granularity dynamics, multiple planning objective trade-offs, and adaptive processes. 3) Data-wise, the field depends heavily on city-specific datasets (e.g., Beijing, NYC), which lack cross-city semantic richness, knowledge, and generalizability and exhibit limited inclusion of human or community feedback. 3) Computation-wise, GenAI models are resource-intensive and prone to instability, especially in adversarial or variational settings; they lack real-time adaptability required for time-sensitive contexts like disaster response, and often remain opaque to end users. 4) Application-wise, even state-of-the-art models (e.g., HRUP) are confined to academic benchmarks, with limited real-world deployment, weak support for human-in-the-loop engagement, and undefined evaluation metrics. These gaps underscore the need to rethink how AI models integrate urban theory, interact with users, and operate across diverse, dynamic planning environments.
III The Vision: Towards Generative AI Urban Planner
The future of urban planning demands systems that are not only data-driven but also context-aware, adaptive, equitable, and collaborative. We envision an AI Urban Planner that integrates human guidance, environmental complexity, and evolving societal priorities to support the full planning life cycle: from ideation and simulation to decision-making and adaptation. Unlike conventional tools, which are rigid and retrospective, the AI Urban Planner will be dynamic, generative, and interactive, capable of co-evolving with cities and communities. We next first propose a principled representation-generation framework for generative urban planning, and then discuss an example application.
III-A A Generic and Principled Generative Urban Planning Framework
@@
III-A1 Representation: Modeling Urban Contexts and Human Needs
@@
III-A2 Generation: Producing Optimal Land-Use and Built Environment Configurations
Given a learned representation, the generative stage synthesizes land-use plans using deep generative models. Several model families are suitable for this task: 1) Variational Autoencoders (VAEs) model a latent variable space that allows for structured sampling and interpolation of land-use patterns. 2) Generative Adversarial Networks (GANs) learn to produce realistic configurations by competing a generator against a discriminator trained to distinguish between real and synthetic plans. 3) Autoregressive Models (e.g., PixelCNN-style) sequentially generate elements of the plan (e.g., grid cells) based on previously generated elements, capturing spatial dependencies. 4) Flow-Based Models apply invertible transformations to map simple distributions into complex urban layouts, supporting exact likelihood estimation. 5) Energy-Based Models define a global energy landscape where desirable configurations have lower energy, allowing generation via gradient-based sampling. These generatie urban planning models can be trained with a combination of generic objectives (e.g., reconstruction loss, likelihood maximization, adversarial loss) and domain-specific constraints tailored to urban planning. These domain-specific constraints balance a diverse mixture of objectives spanning spatial, social, economic, environmental, and governance domains: 1) Spatial objectives focus on land-use compatibility, connectivity, and density optimization to support functional and accessible urban form. 2 )Social objectives emphasize equity, inclusion, community cohesion, and housing affordability to foster livable and just environments. accessible urban form. 3) Economic objectives include maximizing land value, supporting local economies, and minimizing infrastructure costs. 4) Environmental objectives aim to enhance sustainability, reduce emissions, and preserve natural ecosystems. 5) Governance objectives demand compliance with legal frameworks, responsiveness to stakeholder input, and adaptability over time. These objectives often conflict—for example, between maximizing density and preserving green space—highlighting the need for AI systems that can model, navigate, and transparently negotiate such multi-objective trade-offs in planning support tools.
@@ III-A2 Generation: Producing Optimal Land-Use and Built Environment Configurations
III-B Embedding Generative Urban Planning to Real Systems: Flooding Resilience
@@
IV Future Directions and Call to Action
@@ IV-B Differentiate GenAI in Strategic Macro Planning and Scenario Micro Planning
IV-C Integrating Urban Planning Theory via Guided Generation
Generative urban planning need to be guided by urban planning theory. Key theories include: 1) land-use suitability theory that evaluates spatial allocations based on accessibility, environmental constraints, and infrastructure compatibility; 2) central place theory that provides distance-decay models for service distribution and hierarchical urban structures; 3) space syntax theory that quantifies spatial configuration, connectivity, and visibility to model pedestrian flows and urban morphology; 4) resilience and equity planning theory that defines performance metrics related to hazard exposure, service accessibility, and social vulnerability.
These theoretical constructs can be incorporated into generative urban planning models as structural priors, spatial constraints, or optimization objectives. To support theory-guided or conditioned generation, we can employ advanced generative models such as conditional Variational Autoencoders (cVAEs), conditional Generative Adversarial Networks (cGANs), diffusion models with controllable sampling, and reinforcement learning–guided generators, to steer the generative process toward desired urban design plans aligned with planning theory.
IV-D Human-Machine Collaborative Planning
@@
IV-E Leveraging Digital Twins For A Close Simulation, Measurement, Planning Loop
@@
IV-F VLMs and Agentic AI for Urban Planning
To go beyond GenAI like VAEs, GANs, or diffusion models, we believe that VLMs integrate visual and textual modalities and enable more semantically grounded, instruction-following generation. For example, we can align urban design prompts (e.g., “generate a transit-oriented mixed-use layout”) with satellite imagery, GIS data, and street-level visuals, in order to allow planners to interact with models using natural language while preserving spatial fidelity. Agentic AI further models the complex urban planning pipeline by introducing planning, reasoning, task decomposition, and multi-step decision-making abilities. Agent-based urban planners can autonomously explore spatial trade-offs, simulate long-term effects, and iteratively refine plans based on feedback, environmental constraints, or human instructions.
VLMs and agentic systems enable an AI urban planner that is not only generative, but goal-driven, context-aware, and aligned with multimodal planning tasks that are capable of understanding complex instructions, grounding outputs in real-world data, and dynamically adapting to evolving objectives.

### v2

# Towards Urban Planing AI Agent in the Age of Agentic AI

@@ Abstract
Generative AI, large language models, and agentic AI have emerged separately of urban planning. However, the convergence between AI and urban planning presents an interesting opportunity towards AI urban planners. Existing studies conceptualizes urban planning as a generative AI task, where AI synthesizes land-use configurations under geospatial, social, and human-centric constraints and reshape automated urban design. We further identify critical gaps of existing generative urban planning studies: 1) the generative structure has to be predefined with strong assumption: all of adversarial generator-discriminator, forward and inverse diffusion structures, hierarchical zone-POI generative structure are predefined by humans; 2) ignore the power of domain expert developed tools: domain urban planners have developed various tools in the urban planning process guided by urban theory, while existing pure neural networks based generation ignore the power of the tools developed by urban planner practitioners. To address these limitations, we outline a future research direction agentic urban AI planner, calling for a new synthesis of agentic AI and participatory urbanism.
@@ I Introduction and Motivation
II Existing Studies: Urban Planning as Generative AI
@@
II-A Recent Studies in GenAI for Urban Planning
@@
Drawing on the advanced capabilities of generative AI, we have reenvisioned urban planning as a generative AI task, conceptualizing spatial configuration plans as images, matrices, or tensors, with a generative model tasked with producing the optimal tensor representation [1]. The work in [2] frames urban planning as the generation of land-use configurations over spatially attributed graphs augmented with geographic context. These configurations are encoded as multi-channel tensors representing Point of Interest (POI) distributions based on location and function. A graph autoencoder generates regional embeddings, which are fed into a GAN to assess generated plans using metrics such as willingness to pay, diversity of human activities, and social interaction. Despite its novelty, this approach struggles to model spatial hierarchies and exhibits training instability. In [3], land-use generation is conditioned on spatial context and textual planner guidance through a deep conditional variational autoencoder, yielding zone maps and POI layouts informed by contextual factors like geography and mobility. While effective in integrating planning expertise, it does not explicitly address hierarchical urban structures. To overcome these limitations, [4, 1] introduces a hierarchical conditional transformer that first produces zone-level functions before refining them into detailed grid-level plans. Similarly, [5] formulates planning as a two-tier hierarchical reinforcement learning (RL) problem, employing an actor-critic network for zone-level planning and a Deep Q-Network for POI-level design. Likewise, [6] models cities as graphs, utilizing GNN-based RL to make sequential planning decisions.
II-B The Underlying Generative Urban Planning Framework
@@
II-B1 Representation: Modeling Urban Contexts and Human Needs
@@
II-B2 Generation: Producing Optimal Land-Use and Built Environment Configurations
Given a learned representation, the generative stage synthesizes land-use plans using deep generative models, such as, Variational Autoencoders (VAEs), Generative Adversarial Networks (GANs), Autoregressive Models, Flow-Based Models, Energy-Based Models. These generatie urban planning models can be trained with a combination of generic objectives (e.g., reconstruction loss, likelihood maximization, adversarial loss) and domain-specific constraints tailored to urban planning. These domain-specific constraints balance a diverse mixture of objectives spanning spatial, social, economic, environmental, and governance domains: 1) Spatial objectives focus on land-use compatibility, connectivity, and density optimization to support functional and accessible urban form. 2 )Social objectives emphasize equity, inclusion, community cohesion, and housing affordability to foster livable and just environments. accessible urban form. 3) Economic objectives include maximizing land value, supporting local economies, and minimizing infrastructure costs. 4) Environmental objectives aim to enhance sustainability, reduce emissions, and preserve natural ecosystems. 5) Governance objectives demand compliance with legal frameworks, responsiveness to stakeholder input, and adaptability over time. These objectives often conflict, for example, between maximizing density and preserving green space—highlighting the need for AI systems that can model, navigate, and transparently negotiate such multi-objective trade-offs in planning support tools.
@@ II-B2 Generation: Producing Optimal Land-Use and Built Environment Configurations
II-C Embedding Generative Urban Planning to Real Systems: Flooding Resilience
@@
II-D Limitations of Current AI Urban Planner Research
Despite advances in AI for urban planning, current approaches face critical limitations across five foundational dimensions: theory, methodology, data, computing, and applications. 1) Theory-wise, most AI models reduce planning to spatial optimization or land-use generation, thus, ignore core urban concepts such as zoning, resilience, social equity, and participatory governance, and often ignore the political and normative nature of planning decisions—undermining interpretability and legitimacy. 2) Method-wise, models [2, 6, 5, 4, 1, 3, 2] rely on single-step or loosely staged generation conditioned on static inputs that limit creativity and adaptivity, fail to capture multi-granularity dynamics, multiple planning objective trade-offs, and adaptive processes. 3) Data-wise, existing city-specific datasets (e.g., Beijing, NYC) lacks clear evaluation criteria on scoring a good planning and a bad planning. 3) Computation-wise, GenAI models are resource-intensive and prone to instability, especially in adversarial or variational settings; they lacks explainability and transparency due to the use of neural network based architectures, and often remain opaque to end users. 4) Application-wise, existing generative planning methods ignore many domain expert developed tools. Be sure to notice that domain urban planners have developed various tools to assist each step of urban planning process guided by planning theory, while existing pure neural networks based generation ignore the power of the tools developed by urban planner practitioners.
III The Vision: Integrating VLM, Multimodal Reasoning, and Agentic AI for Urban Planning AI Agents
To go beyond GenAI like VAEs, GANs, or diffusion models, we believe that VLMs integrate visual and textual modalities and enable more semantically grounded, instruction-following generation. For example, we can align urban design prompts (e.g., “generate a transit-oriented mixed-use layout”) with satellite imagery, GIS data, and street-level visuals, in order to allow planners to interact with models using natural language while preserving spatial fidelity. Agentic AI further models the complex urban planning pipeline by introducing planning, reasoning, task decomposition, and multi-step decision-making abilities. Agent-based urban planners can autonomously explore spatial trade-offs, simulate long-term effects, and iteratively refine plans based on feedback, environmental constraints, or human instructions. VLMs and agentic systems enable an AI urban planner that is not only generative, but goal-driven, context-aware, and aligned with multimodal planning tasks that are capable of understanding complex instructions, grounding outputs in real-world data, and dynamically adapting to evolving objectives.
[To BE FINISHED]
IV What to Consider in Learning Urban Planning AI Agents
@@ IV-B Differentiate GenAI in Strategic Macro Planning and Scenario Micro Planning
IV-C Human-Machine Collaborative Planning
@@
IV-D Leveraging Digital Twins For A Close Simulation, Measurement, Planning Loop